OpenStudy (aaronandyson):

If charge on one electron is 1.6*10^-19,total charge on 1 mole of electron will be?

OpenStudy (aaronandyson):

@Abhisar

OpenStudy (abhisar):

You tell me, this is very easy.

OpenStudy (abhisar):

How many electrons in 1 mole?

OpenStudy (aaronandyson):

Do I have to multiple the charge,by avogardos number?

OpenStudy (abhisar):

Yes!

OpenStudy (abhisar):

Getting it?

OpenStudy (aaronandyson):

In the question the charge is given as "negative" and yes.

OpenStudy (abhisar):

I know, charge on electron is negative and proton is positive.

OpenStudy (aaronandyson):

so 1.6*6022?

OpenStudy (aaronandyson):

1.6*10^-19 times 6.022*10^23 so it will bw 10^4 cuz -19+23 = 4.

OpenStudy (abhisar):

Correct! \(\sf-1.6 \times 10^{-19} \times 6.022 \times 10^{23}\)

OpenStudy (aaronandyson):

so, 16*6022?

OpenStudy (abhisar):

\(\sf -9.6 \times 10^4\)

OpenStudy (aaronandyson):

How'd you get that?

OpenStudy (abhisar):

Just multiply \(\sf-1.6 \times 10^{-19} \times 6.022 \times 10^{23}\)

OpenStudy (abhisar):

-1.6 x 6.022 = ?

OpenStudy (aaronandyson):

9.635

OpenStudy (abhisar):

Yes, and finally \(\sf 10^{-19} \times 10^{23} = 10^{23-19}=?\)

OpenStudy (aaronandyson):

10^4

OpenStudy (abhisar):

So, the final answer is -9.6 x \(\sf 10^4\)

OpenStudy (abhisar):

Is it clear?

OpenStudy (aaronandyson):

Yes.